# thanks: Corresponding authors:

### v1

@@ G.1 Definition of the effective Hill coefficient
We defined $\hat{w}$ with an extra parameter $h$ to allow a mapping between the thermodynamic model and the Hill function. For the Hill case, we set $h=0$, $\bar{r}_{1}=0$, and $\omega=1$. In our thermodynamic model, $h=1$ and the expression of Eqn. 125 becomes
@@
G.2 Numerical sweep for minimal cooperativity above which there is bistability.
@@
Appendix H Relaxation timescale to equilibrium for the auto-activation system
@@
Figure 33: Relaxation timescales as a function of the initial concentration of gene $A$ ($\bar{A}_{0}$). The parameters of the system are fixed at the following values: $\omega=7.5$, $\bar{r}_{0}=0.1$, $\bar{r}_{1}=1$, $\bar{r}_{2}=20$, and $c=2\cdot 10^{-5}\text{M}$. Timescales are computed using two approaches: a threshold-based method (orange curves) and exponential curve fitting (blue curve). Each orange curve corresponds to a different threshold value $\epsilon$ indicated as percentages in the legend. The threshold timescale $\bar{t}^{*}$ is the time required for $\bar{A}(\bar{t})$ to reach a fraction $\epsilon$ of the total change from initial condition $\bar{A}_{0}$ to steady state $\bar{A}_{\infty}$. The blue curve represents the relaxation time obtained from exponential fits to $\bar{A}(\bar{t})$ trajectories. Vertical lines mark positions of stable (solid lines) and unstable (dashed line) fixed points. The horizontal black line indicates the unity timescale as a reference for comparison.
@@
Appendix I Bistability regimes in the mutual repression circuit
I.1 A necessary condition for bistability
@@
where for convenience we define $p_{1}=p_{\text{act}}(c_{1})$ and $p_{2}=p_{\text{act}}(c_{2})$. By substituting the expression for $\bar{R}_{1}$ of Eqn. 135 into the Eqn. 136 for $\bar{R}_{2}$, Eqn. 136 then becomes
$\bar{R}_{2}=\dfrac{\bar{r}}{1+\dfrac{2p_{1}}{\bar{K}}\dfrac{\bar{r}}{1+2p_{2}\bar{R}_{2}+\omega_{2}(p_{2}\bar{R}_{2})^{2}}+\omega_{1}\left(\dfrac{p_{1}}{\bar{K}}\dfrac{\bar{r}}{1+2p_{2}\bar{R}_{2}+\omega_{2}(p_{2}\bar{R}_{2})^{2}}\right)^{2}}.$ (137)
We can rewrite Eqn. 137 in standard polynomial form, $M(\bar{R}_{2})=0$, where
$M(\bar{R}_{2})=$ $p_{2}^{4}\omega_{2}^{2}\bar{R}_{2}^{5}+p_{2}^{3}\omega_{2}(4-\bar{r}\omega_{2}p_{2})\bar{R}_{2}^{4}$ (138)
@@
$+(1+\frac{2p_{1}\bar{r}}{\bar{K}}+\frac{\omega_{1}p_{1}^{2}\bar{r}^{2}}{\bar{K}^{2}}-4p_{2}r)\bar{R}_{2}-\bar{r}.$
@@
$02+\omega_{2}\left(1+\bar{r}\left(\frac{p_{1}}{\bar{K}}-2p_{2}\right)\right)>$ 0 . (141)
@@
$01+\bar{r}\left(\frac{p_{1}}{\bar{K}}-p_{2}-\frac{\omega_{2}p_{2}}{2}\right)>$ 0 . (142)
@@
The bound stated in Eqn. 150 depends on both $\omega_{2}$ and $\bar{r}$, again, similarly to the auto-activation system, acting together to determine wether bistability can be accessed or not.
@@
From Eqns.  152 and  151 a necessary condition for bistability for $K=1$ is that
@@
I.2 Effective Hill coefficient of the production terms
In the case of auto-activation, analyzing the effective Hill coefficient provided insight into how bistability can arise even in non-cooperative systems—$\omega>1$ is not a necessary condition for bistability, but from our numerical sweeps displayed in Fig. 32, $n_{\text{eff}}>1$ is. Motivated by this, we now examine whether a similar criterion might help explain the restriction of bistability to specific zones of parameter space in mutual repression circuits.
We defined (Eqn. 129, Eqn. 131) and derived an analytical formula (Eqn. 128, Eqn. 127) for the effective Hill coefficient for a general production term
@@ I.2 Effective Hill coefficient of the production terms
in Appendix  G.1. In the mutual repression system, the production terms of interest are the production of $R_{1}$ driven by promoter 1 and regulated by $R_{2}$
@@
With $\bar{r}_{0}=\bar{r}$, $\bar{r}_{1}=0$, and $\bar{r}_{2}=0$, $p_{\text{act}}\equiv p_{\text{act}}(c_{2})$ for $f_{1}(\bar{R}_{2})$ and $p_{\text{act}}\equiv p_{\text{act}}(c_{1})/K$ for $f_{2}(\bar{R}_{1})$; we see that those two production terms fall in to the more general from $w(x)$ of Eqn. 155. We can therefore apply the reasoning and algebra derived in Appendix  G.1. The corresponding effective Hill coefficients are given by
@@
$12\omega_{2}\;>\;\sqrt{1+4\omega_{2}}-$ 1 . (160)
@@
I.3 Effect of cooperativity on the bistability region
@@
Appendix J Separatrix for mutual repression
@@
Appendix K Coherent feed-forward loop response to a step function signal
K.1 Analytical solution for the output $\bar{Z}(\bar{t})$
@@
[…]
In the cases of interest here, $G(t)$ is either $\bar{Y}(\bar{t})$ or $\bar{Z}(\bar{t})$, and $f(t)$ correspondingly is either $f_{Y}(\bar{t})$ or $f_{Z}(\bar{t})$, the activation terms for $Y$ and $Z$, respectively, in Eqns. 167 and  166.
We first solve Eqn. 166, because its dynamics is not coupled to $\bar{Z}(\bar{t})$. Notice that for $\bar{t}>0$, $f_{Y}(\bar{t})$ is constant. Referring to Eqn. 173, we see that $\bar{Y}$ evolves from initial to final state purely exponentially according to the time evolution
@@
with $\mathcal{Y}_{f/i}=p_{\text{act}}^{Y}(c_{Y}^{0})\bar{Y}_{f/i}/\bar{K}_{YZ}$ and $\Delta\mathcal{Y}=\mathcal{Y}_{f}-\mathcal{Y}_{i}$. Given this expression for $\mathcal{Y}(\bar{t})$, we now know the full time dependence of $f_{Z}(\bar{t})$ in Eqn. 167. Next, we can solve for the dynamics of $\bar{Z}$ by substituting $f_{Z}(\bar{t})$ into the general solution Eqn. 173. By evaluating the integral, we find that
[…]
@@
K.2 Derivation and sign of the average delay $\langle\Delta\bar{t}\rangle$
@@
With Eqn. 179, we can analytically evaluate the integral and obtain the following expression for the average delay
$\langle\Delta\bar{t}\rangle=\frac{\Phi(\bar{Z}_{f}-\bar{Z}_{i})^{-1}}{S(1+\omega\mathcal{X}_{f})}\log\Big(\frac{1+\mathcal{X}_{f}+\mathcal{Y}_{i}+\omega\mathcal{X}_{f}\mathcal{Y}_{i}}{S}\Big{missing}).$ (184)
As a reminder, $\langle\Delta\bar{t}\rangle$ signifies the average time difference between the feed-forward loop response and the simple regulation response. The sign of $\langle\Delta\bar{t}\rangle$ indicates whether the feed-forward loop delays ($\langle\Delta\bar{t}\rangle<0$) or accelerates ($\langle\Delta\bar{t}\rangle>0$). From Eqn. 184, we can analytically determine whether the feed-forward loop delays or accelerates by treating the contribution from each component. To begin with, we have $S>0$ and $01+\omega\mathcal{X}_{f}>$ 0 as concentrations are strictly non-negative. For the coherent feed-forward loop, we have $\Phi\geq 0$ as $\bar{r}_{2Z}\geq\bar{r}_{1Z}\geq\bar{r}_{0Z}$ since both $X$ and $Y$ activate $Z$. The term $(\bar{Z}_{f}-\bar{Z}_{i})$ depends on the direction of the step. For an ON step, $(\bar{Z}_{f}-\bar{Z}_{i})>0$ and for an OFF step, $(\bar{Z}_{f}-\bar{Z}_{i})<0$. Finally, the logarithm also depends on the direction of the step. For an ON step, we have
@@
$\implies\log\Big(\frac{1+\mathcal{X}_{f}+\mathcal{Y}_{i}+\omega\mathcal{X}_{f}\mathcal{Y}_{i}}{S}\Big{missing})\leq 0,$ (186)
@@
$\implies\log\Big(\frac{1+\mathcal{X}_{f}+\mathcal{Y}_{i}+\omega\mathcal{X}_{f}\mathcal{Y}_{i}}{S}\Big{missing})\geq 0.$ (188)
Together with other terms, we find $\langle\Delta\bar{t}\rangle\leq 0$ for the OFF step as well.
We have shown that the average time difference $\langle\Delta\bar{t}\rangle$ is negative for both the ON and OFF steps. To complete the analysis, we will further demonstrate that the time difference $\Delta\bar{t}(\bar{Z})$, defined in Sec. IV.1, has the same sign for any concentration $\bar{Z}$. This amounts to saying that the feed-forward trajectory and the simple regulation trajectory never cross each other. We can show this by proving that $\Theta(\bar{t})$ has the same sign for any $\bar{t}>0$. In Eqn. 179, we observe that the time dependence in $\Theta(\bar{t})$ appears in $e^{-\bar{t}}$ and in $Se^{\bar{t}}$ inside the logarithm. $e^{-\bar{t}}>0$ for any $\bar{t}$, thus only the logarithm term might change its sign as time evolves. However, we observe that
@@
K.3 Logic gates analysis in coherent feed-forward loop
@@
$\langle\Delta\bar{t}\rangle$ $=\frac{1}{(\bar{Z}_{f}-\bar{Z}_{i})}\frac{\bar{r}_{1Z}-\bar{r}_{0Z}}{1+\mathcal{X}_{f}+\mathcal{Y}_{f}}\log\Big(\frac{1+\mathcal{X}_{f}+\mathcal{Y}_{i}}{1+\mathcal{X}_{f}+\mathcal{Y}_{f}}\Big{missing}).$ (190)
@@
The ratio between $\langle\Delta\bar{t}\rangle$ for OFF and ON steps is then
$\begin{split}\left|\frac{\langle\Delta\bar{t}\rangle_{\text{OFF}}}{\langle\Delta\bar{t}\rangle_{\text{ON}}}\right|&=\left|\frac{1+\mathcal{X}_{\max}+\mathcal{Y}_{\max}}{1+\mathcal{X}_{\min}+\mathcal{Y}_{\min}}\frac{\log(\frac{1+\mathcal{X}_{\min}+\mathcal{Y}_{\max}}{1+\mathcal{X}_{\min}+\mathcal{Y}_{\min}})}{\log(\frac{1+\mathcal{X}_{\max}+\mathcal{Y}_{\min}}{1+\mathcal{X}_{\max}+\mathcal{Y}_{\max}})}\right|\\[5.0pt] &=\frac{1+\mathcal{X}_{\max}+\mathcal{Y}_{\max}}{1+\mathcal{X}_{\min}+\mathcal{Y}_{\min}}\frac{\log(\frac{1+\mathcal{X}_{\min}+\mathcal{Y}_{\max}}{1+\mathcal{X}_{\min}+\mathcal{Y}_{\min}})}{\log(\frac{1+\mathcal{X}_{\max}+\mathcal{Y}_{\max}}{1+\mathcal{X}_{\max}+\mathcal{Y}_{\min}})}\\[5.0pt] \end{split}$ (191)
@@
$\log\frac{1+\mathcal{X}_{\min}+\mathcal{Y}_{\max}}{1+\mathcal{X}_{\min}+\mathcal{Y}_{\min}}\geq\log\frac{1+\mathcal{X}_{\max}+\mathcal{Y}_{\max}}{1+\mathcal{X}_{\max}+\mathcal{Y}_{\min}}$ (193)
@@
K.4 Analytic solutions of other feed-forward loop networks.
In any feed-forward loop, $X$ regulates $Y$ and $Z$, and $Y$ regulates $Z$. As there are three regulatory interactions, there exist $82^{3}=$ 8 different regulatory logics. For the dynamical equation of $\bar{Y}$, if $X$ activates $Y$, then
@@
There exist 4 different possibilities for how $X$ and $Y$ regulate $Z$. In the main text, we showed the case where $X$ and $Y$ both activate $Z$, and the case where $X$ activates $Z$ but $Y$ represses $Z$. If $X$ represses $Z$, but $Y$ activates $Z$, we have
@@
If both $X$ and $Y$ represses $Z$, then
@@
Using the procedures described in the main text, we can similarly find analytic solutions describing all the other regulatory logics of the feed-forward architectures. Note that for a step function signal in $c_{X}$ and $c_{Y}$, we can always write
@@
Appendix L Functionality condition comparison
In Sec. IV.2, we mentioned that besides the time delay, another helpful criterion for a functional feed-forward loop is the existence of a large difference between the maximum and minimum steady state values of $Z$. Let’s define $\Delta\bar{Z}=|\bar{Z}_{f}-\bar{Z}_{i}|$. We want this to be big so that the dynamical change between low and high concentrations is meaningful. The steady state concentrations $\bar{Z}_{f}$ and $\bar{Z}_{i}$ have expressions given by
@@
Appendix M Technical details in incoherent feed-forward loop
In this appendix, we will return to some technical details regarding pulses in the incoherent feed-forward loop, as presented in Sec. IV.3. To begin with, we discuss the quantity $\langle\Delta\bar{t}\rangle$ in incoherent feed-forward loops. In the coherent feed-forward loop, the definition of $\langle\Delta\bar{t}\rangle$ in Eqn. 36 and Eqn. 37 are equivalent, allowing us to interpret it as the average time delay across concentrations. Here, when $\bar{Z}$ does not exhibit a pulse, there is no difference from the coherent case. However, Eqn. 36 becomes ill-defined when $\bar{Z}$ exhibits a pulse. This is because the feed-forward loop response can reach $\bar{Z}$ in a way that the simple regulation curve cannot. A value can still be computed for $\langle\Delta\bar{t}\rangle$ using Eqn. 37, but it contains information about the relative size of the pulse rather than acceleration. Because it has unit of time, we regard the direct magnitude of the pulse $\bar{Z}_{\max}-\bar{Z}_{f}$ for an increasing $\bar{Z}$ response in the main text.
@@
[…]
@@
Appendix N Effect of $\bar{Y}$ in continuous tuning
@@
In Sec. IV.4, we discussed the scenario where the signal $c_{X}(\bar{t})$ is no longer a step function, but a continuous function in time. We mentioned that the comparison with simple regulation in this case is subtle, as the arbitrary choice of value $\bar{Y}$ affects the shape of simple regulation response. Here, we expand on Fig. 25 where we repeat the numerical integration for three distinct values of $\bar{Y}$, as shown in Fig. 38. For the fast tuning case, different choices of $\bar{Y}$ have no effect on the simple regulation trajectory. This is expected since in the limit case of a step function signal in $c_{X}$, the value of $\bar{Y}$ strictly has no effect on the shape of the trajectory. As the rate of tuning $c_{X}$ slows down, the effect of $\bar{Y}$ becomes more and more pronounced. Specifically, as $\bar{Y}$ increases, the rescaled simple regulation curve “shrinks”. As a result, the magnitude of apparent delay/acceleration between the feed-forward loop and simple regulation response increases. Due to the artificial nature of the choice of $\bar{Y}$ in simple regulation, we cannot make a claim about the magnitude of delay/acceleration in the slow tuning limit. Nevertheless, our qualitative results stand. The choice of $\bar{Y}$ does not change the type of response to a step. For example, even though $\bar{Y}$ affects the magnitude of delay on the OFF step, the feed-forward loop in this case always delays on the OFF step.
Appendix O Code availability

### v2

@@ H.1 Definition of the effective Hill coefficient
We defined $\hat{w}$ with an extra parameter $h$ to allow a mapping between the thermodynamic model and the Hill function. For the Hill case, we set $h=0$, $\bar{r}_{1}=0$, and $\omega=1$. In our thermodynamic model, $h=1$ and the expression of Eqn. 146 becomes
@@
H.2 Numerical sweep for minimal cooperativity above which there is bistability.
@@
Appendix I Relaxation timescale to equilibrium for the auto-activation system
@@
Figure 33: Relaxation timescales as a function of the initial concentration of gene AA ($\bar{A}_{0}$). The parameters of the system are fixed at the following values: $\omega=7.5$, $\bar{r}_{0}=0.1$, $\bar{r}_{1}=1$, $\bar{r}_{2}=20$, and $c=2\cdot 10^{-5}\text{M}$. Timescales are computed using two approaches: a threshold-based method (orange curves) and exponential curve fitting (blue curve). Each orange curve corresponds to a different threshold value $\epsilon$ indicated as percentages in the legend. The threshold timescale $\bar{t}^{*}$ is the time required for $\bar{A}(\bar{t})$ to reach a fraction $\epsilon$ of the total change from initial condition $\bar{A}_{0}$ to steady state $\bar{A}_{\infty}$. The blue curve represents the relaxation time obtained from exponential fits to $\bar{A}(\bar{t})$ trajectories. Vertical lines mark positions of stable (solid lines) and unstable (dashed line) fixed points. The horizontal black line indicates the unity timescale as a reference for comparison.
@@
Appendix J Bistability regimes in the mutual repression circuit
J.1 A necessary condition for bistability
@@
where for convenience we define $p_{1}=p_{\text{act}}(c_{1})$ and $p_{2}=p_{\text{act}}(c_{2})$. Substituting the expression for $\bar{R}_{1}$ of Eqn. 156 into Eqn. 157 for $\bar{R}_{2}$, and rewriting in standard polynomial form, $M(\bar{R}_{2})=0$, we obtain
$M(\bar{R}_{2})=$ $p_{2}^{4}\omega_{2}^{2}\bar{R}_{2}^{5}+p_{2}^{3}\omega_{2}(4-\bar{r}\omega_{2}p_{2})\bar{R}_{2}^{4}$
@@
$+(1+\frac{2p_{1}\bar{r}}{\bar{K}}+\frac{\omega_{1}p_{1}^{2}\bar{r}^{2}}{\bar{K}^{2}}-4p_{2}r)\bar{R}_{2}-\bar{r}.$ (158)
@@
$2+\omega_{2}\left(1+\bar{r}\left(\frac{p_{1}}{\bar{K}}-2p_{2}\right)\right)>0.$ (161)
@@
$1+\bar{r}\left(\frac{p_{1}}{\bar{K}}-p_{2}-\frac{\omega_{2}p_{2}}{2}\right)>0.$ (162)
@@
The bound stated in Eqn. 170 depends on both $\omega_{2}$ and $\bar{r}$, again, similarly to the auto-activation system, acting together to determine wether bistability can be accessed or not.
@@
From Eqns. 172 and 171 a necessary condition for bistability for $K=1$ is that
@@
J.2 Effective Hill coefficient of the production terms
[…]
We defined (Eqn. 150, Eqn. 152) and derived an analytical formula (Eqn. 149, Eqn. 148) for the effective Hill coefficient for a general production term
@@
in Appendix H.1. In the mutual repression system, the production terms of interest are the production of $R_{1}$ driven by promoter 1 and regulated by $R_{2}$
@@
With $\bar{r}_{0}=\bar{r}$, $\bar{r}_{1}=0$, and $\bar{r}_{2}=0$, $p_{\text{act}}\equiv p_{\text{act}}(c_{2})$ for $f_{1}(\bar{R}_{2})$ and $p_{\text{act}}\equiv p_{\text{act}}(c_{1})/K$ for $f_{2}(\bar{R}_{1})$; we see that those two production terms fall in to the more general from $w(x)$ of Eqn. 175. We can therefore apply the reasoning and algebra derived in Appendix H.1. The corresponding effective Hill coefficients are given by
@@
$2\omega_{2}\;>\;\sqrt{1+4\omega_{2}}-1.$ (180)
@@
J.3 Effect of cooperativity on the bistability region
@@
Appendix K Separatrix for mutual repression
@@
Appendix L Coherent feed-forward loop response to a step function signal
L.1 Analytical solution for the output $\bar{Z}(\bar{t})$
@@
$G(t)=e^{-t}\Big(G(0)+\int_{0}^{t}e^{t^{\prime}}f(t^{\prime})dt^{\prime}\Big).$ (193)
In the cases of interest here, $G(t)$ is either $\bar{Y}(\bar{t})$ or $\bar{Z}(\bar{t})$, and $f(t)$ correspondingly is either $f_{Y}(\bar{t})$ or $f_{Z}(\bar{t})$, the activation terms for $Y$ and $Z$, respectively, in Eqns. 187 and 186.
We first solve Eqn. 186, because its dynamics is not coupled to $\bar{Z}(\bar{t})$. Notice that for $\bar{t}>0$, $f_{Y}(\bar{t})$ is constant. Referring to Eqn. 193, we see that $\bar{Y}$ evolves from initial to final state purely exponentially according to the time evolution
@@
with $\mathcal{Y}_{f/i}=p_{\text{act}}^{Y}(c_{Y}^{0})\bar{Y}_{f/i}/\bar{K}_{YZ}$ and $\Delta\mathcal{Y}=\mathcal{Y}_{f}-\mathcal{Y}_{i}$. Given this expression for $\mathcal{Y}(\bar{t})$, we now know the full time dependence of $f_{Z}(\bar{t})$ in Eqn. 187. Next, we can solve for the dynamics of $\bar{Z}$ by substituting $f_{Z}(\bar{t})$ into the general solution Eqn. 193. By evaluating the integral, we find that
$\begin{split}\bar{Z}(\bar{t})&=\bar{Z}_{i}e^{-\bar{t}}+\Big(e^{-\bar{t}}\int_{0}^{\bar{t}}e^{t^{\prime}}f_{Z}(t^{\prime})dt^{\prime}\Big)\\[5.0pt] &=\bar{Z}_{i}e^{-\bar{t}}+\Big(\bar{Z}_{f}(1-e^{-\bar{t}})+\Theta(\bar{t})\Big)\\[5.0pt] &=\bar{Z}_{\rm simple}(\bar{t})+\Theta(\bar{t}),\end{split}$ (198)
@@
L.2 Derivation and sign of the average delay $\langle\Delta\bar{t}\rangle$
@@
$\langle\Delta\bar{t}\rangle=\>\frac{1}{\bar{Z}_{f}-\bar{Z}_{i}}\int_{0}^{\infty}\Theta(\bar{t})d\bar{t}.$ (203)
With Eqn. 199, we can analytically evaluate the integral and obtain the following expression for the average delay
@@
As a reminder, $\langle\Delta\bar{t}\rangle$ signifies the average time difference between the feed-forward loop response and the simple regulation response. The sign of $\langle\Delta\bar{t}\rangle$ indicates whether the feed-forward loop delays ($\langle\Delta\bar{t}\rangle<0$) or accelerates ($\langle\Delta\bar{t}\rangle>0$). From Eqn. 204, we can analytically determine whether the feed-forward loop delays or accelerates by treating the contribution from each component. To begin with, we have $S>0$ and $1+\omega\mathcal{X}_{f}>0$ as concentrations are strictly non-negative. For the coherent feed-forward loop, we have $\Phi\geq 0$ as $\bar{r}_{2Z}\geq\bar{r}_{1Z}\geq\bar{r}_{0Z}$ since both $X$ and $Y$ activate $Z$. The term $(\bar{Z}_{f}-\bar{Z}_{i})$ depends on the direction of the step. For an ON step, $(\bar{Z}_{f}-\bar{Z}_{i})>0$ and for an OFF step, $(\bar{Z}_{f}-\bar{Z}_{i})<0$. Finally, the logarithm also depends on the direction of the step. For an ON step, we have
@@
$\implies\log\Big(\frac{1+\mathcal{X}_{f}+\mathcal{Y}_{i}+\omega\mathcal{X}_{f}\mathcal{Y}_{i}}{S}\Big.)\leq 0,$ (206)
@@
$\implies\log\Big(\frac{1+\mathcal{X}_{f}+\mathcal{Y}_{i}+\omega\mathcal{X}_{f}\mathcal{Y}_{i}}{S}\Big.)\geq 0,$ (208)
and we thus find that $\langle\Delta\bar{t}\rangle\leq 0$ for the OFF step as well.
We have shown that the average time difference $\langle\Delta\bar{t}\rangle$ is negative for both the ON and OFF steps. To complete the analysis, we will further demonstrate that the time difference $\Delta\bar{t}(\bar{Z})$, defined in Section IV.1, has the same sign for any concentration $\bar{Z}$. This amounts to saying that the feed-forward trajectory and the simple regulation trajectory never cross each other. We can show this by proving that $\Theta(\bar{t})$ has the same sign for any $\bar{t}>0$. In Eqn. 199, we observe that the time dependence in $\Theta(\bar{t})$ appears in $e^{-\bar{t}}$ and in $Se^{\bar{t}}$ inside the logarithm. $e^{-\bar{t}}>0$ for any $\bar{t}$, thus only the logarithm term might change its sign as time evolves. However, we observe that
@@
L.3 Logic gate analysis in coherent feed-forward loop
@@
[…]
@@
[…]
@@
$\log\frac{1+\mathcal{X}_{\min}+\mathcal{Y}_{\max}}{1+\mathcal{X}_{\min}+\mathcal{Y}_{\min}}\geq\log\frac{1+\mathcal{X}_{\max}+\mathcal{Y}_{\max}}{1+\mathcal{X}_{\max}+\mathcal{Y}_{\min}}.$ (213)
@@
L.4 Analytic solutions of other feed-forward loop networks.
In any feed-forward loop, $X$ regulates $Y$ and $Z$, and $Y$ regulates $Z$. As there are three regulatory interactions, there exist $2^{3}=8$ different regulatory logics. For the dynamical equation of $\bar{Y}$, if $X$ activates $Y$, then
@@
There exist 4 different possibilities for how $X$ and $Y$ regulate $Z$. In Section IV, we showed the case where $X$ and $Y$ both activate $Z$, and the case where $X$ activates but $Y$ represses $Z$. If $X$ represses and $Y$ activates $Z$, we have
@@
If both $X$ and $Y$ repress $Z$, then
@@
Using the procedures described in the main text, we can similarly find analytic solutions describing all the other regulatory logics of the feed-forward architectures. Note that for a step function signal in $c_{X}$ and $c_{Y}$,
@@
Appendix M Functionality condition comparison
In Section IV.2, we mentioned that besides the time delay, another helpful criterion for a functional feed-forward loop is the existence of a large difference between the maximum and minimum steady state values of $Z$. Let’s define $\Delta\bar{Z}=|\bar{Z}_{f}-\bar{Z}_{i}|$. We want this to be big so that the dynamical change between low and high concentrations is meaningful. The steady state concentrations $\bar{Z}_{f}$ and $\bar{Z}_{i}$ have expressions given by
@@
Appendix N Technical details in incoherent feed-forward loop
In this appendix, we will return to some technical details regarding pulses in the incoherent feed-forward loop, as presented in Section IV.3. To begin with, we discuss the quantity $\langle\Delta\bar{t}\rangle$ in incoherent feed-forward loops. In the coherent feed-forward loop, the definition of $\langle\Delta\bar{t}\rangle$ in Eqn. 36 and Eqn. 37 are equivalent, allowing us to interpret it as the average time delay across concentrations. Here, when $\bar{Z}$ does not exhibit a pulse, there is no difference from the coherent case. However, Eqn. 36 becomes ill-defined when $\bar{Z}$ exhibits a pulse. This is because the feed-forward loop response can reach $\bar{Z}$ in a way that the simple regulation curve cannot. A value can still be computed for $\langle\Delta\bar{t}\rangle$ using Eqn. 37, but it contains information about the relative size of the pulse rather than acceleration. Because it has unit of time, we regard the direct magnitude of the pulse $\bar{Z}_{\max}-\bar{Z}_{f}$ for an increasing $\bar{Z}$ response in the main text.
@@
$\begin{split}\langle\Delta\bar{t}\rangle_{\max}&=\frac{1}{\bar{Z}_{f}-\bar{Z}_{i}}\int_{0}^{\infty}d\bar{t}\left(\bar{Z}_{\text{step}}(\bar{t})-\bar{Z}_{\text{simple}}(\bar{t})\right)\\ &=\frac{1}{\bar{Z}_{f}-\bar{Z}_{i}}\int_{0}^{\infty}d\bar{t}\left(\bar{Z}_{f}-\bar{Z}_{i}e^{-\bar{t}}-\bar{Z}_{f}(1-e^{-\bar{t}})\right)\\ &=\frac{1}{\bar{Z}_{f}-\bar{Z}_{i}}\int_{0}^{\infty}d\bar{t}(\bar{Z}_{f}-\bar{Z}_{i})e^{-\bar{t}}\\ &=\frac{\bar{Z}_{f}-\bar{Z}_{i}}{\bar{Z}_{f}-\bar{Z}_{i}}=1.\end{split}$ (230)
@@
Appendix O Effect of $\bar{Y}$ in continuous tuning
@@
In Section IV.4, we discussed the scenario where the signal $c_{X}(\bar{t})$ is no longer a step function, but a continuous function in time. We mentioned that the comparison with simple regulation in this case is subtle, as the arbitrary choice of value $\bar{Y}$ affects the shape of simple regulation response. Here, we expand on Fig. 25 where we repeat the numerical integration for three distinct values of $\bar{Y}$, as shown in Fig. 38. For the fast tuning case, different choices of $\bar{Y}$ have no effect on the simple regulation trajectory. This is expected since in the limit case of a step function signal in $c_{X}$, the value of $\bar{Y}$ strictly has no effect on the shape of the trajectory. As the rate of tuning $c_{X}$ slows down, the effect of $\bar{Y}$ becomes more and more pronounced. Specifically, as $\bar{Y}$ increases, the rescaled simple regulation curve “shrinks”. As a result, the magnitude of apparent delay/acceleration between the feed-forward loop and simple regulation response increases. Due to the artificial nature of the choice of $\bar{Y}$ in simple regulation, we cannot make a claim about the magnitude of delay/acceleration in the slow tuning limit. Nevertheless, our qualitative results stand. The choice of $\bar{Y}$ does not change the type of response to a step. For example, even though $\bar{Y}$ affects the magnitude of delay on the OFF step, the feed-forward loop in this case always delays on the OFF step.
Appendix P Code availability